OpenStudy (anonymous):

Which of (3, 5), (4, 6), (5, 7), and (6, 8) are solutions to y = x + 2? (1 point)(3, 5) and (4, 6) (4, 6) and (5, 7) none all

OpenStudy (aripotta):

-_- i tried helping you. it's simple. i'm not going to tell you the answer unless you actually work for it

OpenStudy (anonymous):

So then dont tell me thee answer and go elsewhere

OpenStudy (aripotta):

no one's going to just tell you

OpenStudy (anonymous):

Actually that's what everyone does ,

OpenStudy (anonymous):

they are not suppose to

OpenStudy (aripotta):

points on a plane are set up as (x,y) right? so if you substitute the x's and y's of the points into the equation, you can see if they're true. if they're true, then it's a solution

OpenStudy (aripotta):

does 4+2=6?

OpenStudy (aripotta):

does 5+2=7?

OpenStudy (aripotta):

does 6+2=8?

OpenStudy (farmdawgnation):

Sorry, @BeedalButt this isn't a site where you can expect users to just give you the answers. If that's been your experience I apologize - we fell short of our standard in those situations. Please do work with people who are trying to help.

OpenStudy (aripotta):

you're making it difficult when it doesn't need to be. i'm trying to help. i practically just gave you the answer